OpenStudy (anonymous):

which ratio forms a proportion with 3/21 ? 2/35 5/49 2/14 7/35

OpenStudy (anonymous):

3/21 = 1/7 1/7 = 2/14 = 3/21 = 4/28 ... so which one

OpenStudy (anonymous):

uhh 2/14? -_-

OpenStudy (anonymous):

$5